# Entanglement suppression and quantum scars in a three-oscillator gravitational analogue

### v1

@@ Appendix D Log Negativity in the weak coupling limit
$\Delta=-\sin{(2\omega\lambda_{p}t)}R\begin{pmatrix}\mathbb{I}_{2}&\mathbb{I}_{2}\\ \mathbb{I}_{2}&\mathbb{I}_{2}\end{pmatrix}$ )
@@
$\tilde{\nu}_{1+}$$\approx$ $\sqrt{1+\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70a)
$\tilde{\nu}_{1-}$$\approx$ $-\sqrt{1+\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70b)
$\tilde{\nu}_{2+}$$\approx$ $\sqrt{1-\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70c)
$\tilde{\nu}_{2-}$$\approx$ $-\sqrt{1-\lambda_{p}\sin{(2\omega\lambda_{p}t)}}$ (70d)
@@
$\mathcal{E}_{\mathcal{N}}(t)\approx\frac{\lambda_{p}|\sin{(2\sqrt{2}gt)}|}{2\ln{(2)}}+\mathcal{O}({\lambda_{p}^{2}})$ ) (71)
@@ Appendix E Plots for Vacuum Initial state
In Fig. (3) we provide the contour maps for $\langle\mathcal{E}_{\mathcal{N}}\rangle$, and $\langle\mathcal{F}\rangle$ for the case where the initial state of the system is vacuum. For the case of initial squeezed states, the plots were provided in Fig. (2) in the main text [53]. In Fig. (4) above we provide the countour plots for quantum discord $\langle\mathcal{D}\rangle$ both for initial squeezed state and initial vacuum state. Note that the behaviour of quantum discord resembles that of logarithmic negativity.

### v3

@@ Appendix D Log Negativity in the weak coupling limit
[…]
@@
[…]
@@
[…]
@@ Appendix E Plots for Vacuum Initial state
In Fig. (4) we provide the contour maps for $\langle\mathcal{E}_{\mathcal{N}}\rangle$, and $\langle\mathcal{F}\rangle$ for the case where the initial state of the system is vacuum. For the case of initial squeezed states, the plots were provided in Fig. 2. In Fig. 5 above we provide the countour plots for quantum discord $\langle\mathcal{D}\rangle$ both for initial squeezed state and initial vacuum state. Note that the behaviour of quantum discord resembles that of logarithmic negativity.
Appendix F Exact Solution for the HMR Limit
In this section, we present the exact solution for the heavy mediator regime (HMR), characterized by the condition $\omega_{B}=\lambda_{B}$. In this limit, the eigenmode frequencies simplify to $k_{1}=k_{3}=\sqrt{\omega_{A}^{2}-\lambda_{A}^{2}}$ and $k_{2}=0$, with the vanishing of $k_{2}$ indicating a frozen mode. Following the same procedure outlined in Appendix. B, and substituting $\omega_{B}=\lambda_{B}$, the relations between the Bogoliubov coefficients $\alpha_{ij}$ and $\beta_{ij}$ reduce to:
$\beta_{11}$ $=\frac{\alpha_{11}(k_{1}-\omega_{A}+\lambda_{A})}{(k_{1}+\omega_{A}-\lambda_{A})}$ (75)
$\beta_{12}$ $=\alpha_{12}$ (76)
$\beta_{13}$ $=\frac{\alpha_{13}(k_{1}-\omega_{A}+\lambda_{A})}{(k_{1}+\omega_{A}-\lambda_{A})}$ (77)
$\beta_{21}$ $=-\alpha_{21}$ (78)
$\beta_{23}$ $=-\alpha_{23}$ (79)
$\beta_{31}$ $=\frac{\alpha_{31}(k_{3}-\omega_{A}+\lambda_{A})}{(k_{3}+\omega_{A}-\lambda_{A})}$ (80)
$\beta_{32}$ $=\alpha_{32}$ (81)
$\beta_{33}$ $=\frac{\alpha_{33}(k_{3}-\omega_{A}+\lambda_{A})}{(k_{3}+\omega_{A}-\lambda_{A})}$ (82)
While this method does not yield a direct relation for $\beta_{22}$, we may introduce the parametrization $\beta_{22}=\chi_{2}\alpha_{22}$ with $|\chi_{2}|\geq 1$ without loss of generality. The full set of Bogoliubov coefficients is summarized in Table 2, where $\chi_{1}$ is an arbitrary real parameter. Interestingly, in this limit the only constraint arising from the positivity of the eigenfrequencies is $\omega_{A}\geq\lambda_{A}$. This permits arbitrarily large values for the subsystem-mediator couplings $g_{A}$ and $g_{B}$. However, despite this strong coupling, the structure of the Bogoliubov transformation in this regime prohibits any entanglement transfer between the subsystems, effectively rendering the mediator dynamically inert. This can be verified analytically by computing the symplectic eigenvalues of the partial transposed reduced covariance matrix, which turn out to be
$\tilde{\nu}_{1}$ $=1$
[…]
Since $\omega_{A}>\lambda_{A}$, both symplectic eigenvalues are always positive and greater than 1.
Figure 6: Normalized time-averaged dynamical fidelity susceptibility vs system parameters. Left: vacuum initial state; right: squeezed initial state with $r_{1}=r_{3}=0.5$, $r_{2}=0.2$, $\theta_{1}=\theta_{2}=\theta_{3}=0$. Fixed parameters: $\omega_{A}=5$, $\lambda_{A}=0.01\,\omega_{A}$, $g_{A}=g_{B}=0.5$.
Appendix G Dynamical Fidelity Susceptibility
The time-averaged dynamical fidelity susceptibility $\langle\chi_{F}\rangle$ provides a potential method for identifying the mediator’s regime in BMV-type experiments. In this section, we present the results for the variation of $\langle\chi_{F}\rangle$ with $\omega_{B}$ in the case $\lambda_{A}=0.01\,\omega_{A}$. As shown in Fig. 6, the qualitative features closely resemble those observed for $\lambda_{A}=0.5\,\omega_{A}$. In particular, the stability of $\langle\chi_{F}\rangle$ in the HMR limit under variations in $\omega_{B}$ near $\omega_{A}$ suggests that this behavior can serve as a reliable diagnostic for characterizing the nature of the mediator.
Table 2: Expressions for the Bogoliubov coefficients $\alpha_{ij}$ and $\beta_{ij}$ in the HMR limit.
$j=1$ $j=2$ $j=3$
$\alpha_{1j}$ $\dfrac{\chi_{1}(k_{1}+\omega_{A}-\lambda_{A})}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$ $\dfrac{k_{1}(g_{B}+\chi_{1}g_{A})}{(\omega_{A}+\lambda_{A})\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$ $\dfrac{(k_{1}+\omega_{A}-\lambda_{A})}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\alpha_{2j}$ $\dfrac{-(\chi_{2}-1)g_{A}}{(\omega_{A}+\lambda_{A})\sqrt{\chi_{2}^{2}-1}}$ $\dfrac{\chi_{2}}{\sqrt{\chi_{2}^{2}-1}}$ $\dfrac{-(\chi_{2}-1)g_{B}}{(\omega_{A}+\lambda_{A})\sqrt{\chi_{2}^{2}-1}}$
$\alpha_{3j}$ $\dfrac{(k_{1}+\omega_{A}-\lambda_{A})}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ $\dfrac{k_{1}(g_{A}-\chi_{1}g_{B})}{2(\omega_{A}+\lambda_{A})\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$ $\dfrac{(k_{1}+\omega_{A}-\lambda_{A})\chi_{1}}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$
$\beta_{1j}$ $\dfrac{\chi_{1}(k_{1}-\omega_{A}+\lambda_{A})}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$ $\dfrac{k_{1}(g_{B}+\chi_{1}g_{A})}{(\omega_{A}+\lambda_{A})\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$ $\dfrac{(k_{1}-\omega_{A}+\lambda_{A})}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$
$\beta_{2j}$ $\dfrac{(\chi_{2}-1)g_{A}}{(\omega_{A}+\lambda_{A})\sqrt{\chi_{2}^{2}-1}}$ $\dfrac{1}{\sqrt{\chi_{2}^{2}-1}}$ $\dfrac{(\chi_{2}-1)g_{B}}{(\omega_{A}+\lambda_{A})\sqrt{\chi_{2}^{2}-1}}$
$\beta_{3j}$ $\dfrac{(k_{1}-\omega_{A}+\lambda_{A})}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{1}-\lambda_{1})}}$ $\dfrac{k_{1}(g_{A}-\chi_{1}g_{B})}{2(\omega_{A}+\lambda_{A})\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$ $\dfrac{(k_{1}-\omega_{A}+\lambda_{A})\chi_{1}}{2\sqrt{1+\chi_{1}^{2}}\sqrt{k_{1}(\omega_{A}-\lambda_{A})}}$